the length of a rectangle is 5 less than 2 times its width, if the perimeter is 38. what are the dimensions
Answer by nathany1280(5)

You know the length of the rectangle, 5 less than 2 times its width. That would be 2x-5. You don't know the width, so it's x. The total or perimeter is 38. and the formula is 2l+2w. So the equation is 2(2x-5)+2x=38 Then 4x-10+2x=38. then 6x-10=38. Bring the 10 over and would be 6x=48. x=8.
You are not done yet! You have to plug in 8 into 2x-5 for the length. The width you know is 8. The length equation is 2(8)-5. 16-5. then it the length would equal 11.

Length equal 11. Width equal 8.